A screw jack of length 12cm is used to raise a car through a vertical height of 25cm by turning a Tommy bar through 50 revolution

calculate the approximate of the screw jack

(π=3-14)


1
Expert's answer
2021-02-22T07:13:43-0500

The pitch of the screw jack is


"p=\frac{25}{50}=0.5\text{ cm}."

The approximate velocity ratio of the screw jack is


"v=\frac{2\pi D}{p}=\frac{2\pi ·12}{0.5}=151."